A 4-stage pipeline has the stage delays as 150, 120, 160 and 140


nanoseconds respectively. Registers that are used between the stages have a


delay of 5 nanoseconds each.


Question: Assuming constant clock rate, calculate the total time taken to


process 1000 data items on this pipeline.

1
Expert's answer
2022-03-18T02:26:33-0400

Delay between each stage is 5 ns.


Total delay for 1st data item = 165"\times"4

               = 660 

For 1000 data items, first data will take 660 ns to complete and rest 

999 data will take max of all the stages that is 160 ns + 5 ns register delay


Total Delay = 660 + 999"\times"165 ns


=165.5 microsecond.